The length of a rectangular playground is 7 ft more than five times the width. The perimeter is 254 ft. What are the dimensions of rectangle

Answer by checkley77(12844)

L=5W+7
P=2L+2W
254=2(5W+7)+2W
254=10W+14+2W
254=12W+14
12W=254-14
12W=240
W=240/12
W=20 FT.
L=5*20+7
L=100+7=107 ANS FOR THE LENCTH.
PROOF:
254=2*107+2*20
254=214+40
254=254